OpenStudy (anonymous):

What is the area of a square with a side length of 6 ft? i don't know how to do this but if i guessed from the options would i square 6 and pick 36?

OpenStudy (anonymous):

6x6 = 36

OpenStudy (angela210793):

A=x^2=6^2=6*6=36

OpenStudy (angela210793):

A of the square =side^2

OpenStudy (anonymous):

thanks both of you sorry asking so many questions even if there right trying to make sure i know what i need to study for this exam

OpenStudy (angela210793):

yw :)